# SimBench: Benchmarking the Ability of Large Language Models to Simulate Human Behaviors

### v3

@@ 6 Conclusion
Using this benchmark, we provide the first systematic analysis of this capability, showing that even SOTA LLMs have limited simulation ability, performance scales log-linearly with model size, and there exists a fundamental tradeoff between standard alignment and simulating diverse human opinions. We further reveal that models struggle more when simulating specific demographic groups. We also show that strong simulation ability correlates with deep, knowledge-intensive reasoning. While significant progress is needed, SimBench makes this progress measurable, providing an open foundation to accelerate the development of more faithful LLM simulators.
7 Ethics Statement
SimBench’s primary purpose is to benchmark LLMs’ ability to simulate human behavior. While advancements in LLM simulation capabilities can support helpful applications such as pre-testing policies, these do not come without risks of misrepresentation and dual use.
7.1 Responsible Use and Acknowledgment of Limitations
@@
7.2 Data Provenance and Transformative Use
@@
7.3 Scope of Representation and Intersectional Analysis
@@ 7.4 Conclusion
8 Reproducibility Statement
To ensure the reproducibility of our findings and to facilitate future research, we make all components of our work publicly available. The complete SimBench benchmark, including all 20 processed datasets, our curated SimBenchPop and SimBenchGrouped splits, and detailed data cards, will be released on the Hugging Face Hub. Our codebase will be available in an open-source GitHub repository. We provide detailed descriptions of our experimental setup, including the exact prompts used for both base and instruction-tuned models, in Appendix D. We further provide an empirical validation of our elicitation methodology choice in Appendix E.
@@ Appendix A Limitations
The challenge of benchmarking human behavior lies in its vastness. No single study, nor dozens, could ever capture its full complexity. We operationalize this challenge by focusing on a core set of fundamental cognitive and social tasks widely studied in the behavioral sciences: decision-making, self-assessment, judgment, and problem-solving. This focus necessarily means that SimBench does not capture the full complexity of human interaction, such as embodied or multi-turn social dynamics. As the first large-scale benchmark for group-level simulation, SimBench provides the essential infrastructure to establish robust baselines, uncover fundamental properties like scaling laws, and map the frontier for future work on more complex, interactive simulations.
@@ Appendix C SimBenchPop and SimBenchGrouped Sampling Details
For SimBenchGrouped, we focus only on five large-scale survey datasets with rich demographic information and sufficient sample sizes: OpinionQA, ESS, Afrobarometer, ISSP, and LatinoBarometro. Our sampling approach prioritizes questions showing meaningful demographic variation. For each dataset, we identify available grouping variables (e.g., age, gender, country) with sufficient group sizes to form meaningful response distributions. We calculate the variance of responses across demographic groups for each question and rank questions by their variance scores, prioritizing those showing the strongest demographic differences. We select questions that exhibit significant variation across demographic groups to ensure the benchmark captures meaningful differences in responses. For each selected question, we create multiple test cases by pairing it with different values of the grouping variables (e.g., age = “18-29”, age = “30-49”). This process results in 6,343 test cases that specifically measure LLMs’ ability to simulate responses from narrower participant groups based on specified demographic characteristics. Table 4 provides a summary of the sampling process across all datasets, showing the minimum group size thresholds and the number of test cases in each benchmark split.
Table 4: Dataset Sampling Summary; NaN refers to dataset that is only available in aggregated form and no grouping size is known.
@@ Appendix E Validation of Elicitation Method
Figure 5 compares the SimBench scores for several instruction-tuned models using both methods. The results are unequivocal: using verbalized distributions (teal dots) dramatically and consistently outperforms direct token probabilities (orange dots) for every instruction-tuned model tested. In many cases, using token probabilities results in scores far below zero, indicating that the model’s raw logits are poorly calibrated for this task after instruction-tuning. In contrast, base models (black bars) perform reasonably well with token probabilities, as they are not subject to the same post-training shifts.
@@ Appendix G Full SimBench Results (RQ1)
We show the SimBench scores for all the 45 models we evaluate in Table 7. We show the scaling law plots for all models in Figure 6.
@@ H.2 Within-Family Performance Analysis
To investigate the impact of model size (RQ2), we conducted one-way ANOVA followed by post-hoc Tukey HSD tests for pairwise comparisons within model families. The results, presented in Table 9, show that for most families, increases in model size lead to statistically significant improvements in simulation ability, supporting the scaling laws observed in Figure 2.
@@
C(dataset_name)[T.LatinoBarometro] -0.0368 0.0019 -18.9530 0.0000 -0.0406 -0.0330
@@ Conclusion
These results provide strong evidence for inconsistent mediation and resolve a key paradox in our findings. While our analysis in §4.4 shows that instruction-tuning harms simulation fidelity on high-entropy questions, our main leaderboard (Table 4.1) shows that the best overall simulators are instruction-tuned. This mediation analysis explains why: the total effect of instruction tuning is the net outcome of two larger, opposing forces. First, a direct positive effect (+6.46 points) on capability, likely from improved instruction- and persona-following. Second, a smaller but significant indirect negative effect (-1.74 points) caused by entropy suppression. The net positive effect (+4.72 points) demonstrates that, on average, the direct benefits of alignment currently outweigh the harm from reduced distributional diversity. Future work on creating SOTA simulators should therefore focus on developing hybrid or “distribution-preserving” alignment methods that retain the direct benefits of instruction-tuning while mitigating its harmful, entropy-reducing side effects.
@@
This qualitative divergence is key. The two methods are not just different in degree, but in kind. Instruction-tuning boosts performance by sharpening a model’s ability to follow prompt instructions and converge on a consensus answer. Specialist tuning boosts performance by aligning the model’s internal representations more closely with the patterns of human choice. Because they target different mechanisms, their benefits are not mutually exclusive. This suggests that perhaps future gains in LLM simulation will come from hybrid approaches that synthesize both paradigms, creating models that are both generally capable and foundationally aligned with the nuances of human behavior.
@@
Questions: 113, with an average of 518 responses per question.
@@
Participant grouping variables: None. Default System Prompt:
@@
Participants: [roughly 10,000] paid participants from a representative sample of the US populace.
@@
Description: This dataset contains three moral decision-making scenarios, which a large number of participants were asked to provide subjective choices for. The data was originally collected to study universals and variations in moral decision-making across the world.
@@
Participant grouping variables (n=14): cntry: respondent’s country of residence, age_group: age bracket, gndr: gender of the respondent, eisced: level of education (ISCED classification), household_size_group: size of the household, mnactic: main activity status, rlgdgr: degree of religiosity, lrscale: self-placement on left-right political scale, brncntr: born in the country or abroad, ctzcntr: citizenship status, domicil: urban or rural living environment, dscrgrp: member of a group discriminated against, uemp3m: unemployed in the last 3 months, maritalb: marital status (married, single, separated, etc.)
@@
Participant grouping variables (n=11): country: respondent’s country, gender: male or female, education: education level, age_group: age bracket, religion: stated religion, urban_rural: area of living, employment: job situation, bank_account: whether respondent has a bank account, ethnic_group: respondent’s ethnicity, subjective_income: how often to go without cash income, discuss_politics: how often to discuss politics,
@@
Participant grouping variables (n=18): age_group: age bracket, gender_cat: male or female or other, race_cat: respondent’s race, engnat_cat: is (not) English native, hand_cat: right/left/both-handed, education_cat: level of education, urban_cat: type of urban area, religion_cat: stated religion, orientation_cat: sexual orientation, voted: did (not) vote at last elections, married: never/currently/previously, familysize: number of people belonging to the family, TIPI_E_Group: extraversion level based on TIPI score, TIPI_A_Group: agreeableness level based on TIPI score, TIPI_C_Group: conscientiousness level based on TIPI score, TIPI_ES_Group: emotional stability level based on TIPI score, TIPI_O_Group: openness-to-experience level based on TIPI score. household_income: income sufficiency, work_status: job situation, religion: stated religion, nr_of_persons_in_household: 1-7+, marital_status respondent’s legal relationship status, domicil: type of urban area,
@@
Participant grouping variables (n=11): country: respondent’s country, age: age bracket, gender: male or female, years_of_education: 1-10+, household_income: income sufficiency, work_status: job situation, religion: stated religion, nr_of_persons_in_household: 1-7+, marital_status respondent’s legal relationship status, domicil: type of urban area, topbot: self-asessed social class
@@
Looking at the list below, who do you think should have the greatest responsibility for reducing differences in income between people with high incomes and people with low incomes? Options: (A): Cant choose (B): Private companies (C): Government (D): Trade unions (E): High-income individuals themselves (F): Low-income individuals themselves (G): Income differences do not need to be reduced
@@
Participants: In total, 19,205 interviews were applied in 17 countries. Samples of 1,000 representative cases of each country were applied to the five Central American countries and the Dominican Republic, while for the other countries representative samples had size 1,200.
@@
Participant grouping variables (n=1): country
@@
Description: This dataset comprises 990 multi-turn adversarial conversations between humans and a dialogue model, accompanied by raters’ judgement on the safety of last model response. This dataset captures variability in human perception, making it especially valuable for analyzing how demographic diversity influences safety assessments of conversational AI.
Questions: 27,748 (990 turn annotated in multiple dimensions), with an average of 72 responses per question.
@@
Participant grouping variables (n=4): Country: country of origin, Age_Group: age bracket of the respondent, Gender: gender of the respondent, Gender: highest level of education completed
@@
Participant grouping variables (n=1): UserCountry3: participant country,
@@
Description: This dataset includes judgements about individuals’ perception of science, its role in society and politics, attitudes toward climate change, and science communication behaviors.
Questions: 97, with an average of 69.234 responses per question.
@@
Prior research has primarily relied on first token probabilities to obtain survey answers from LLMs (Santurkar et al., 2023; Dominguez-Olmedo et al., 2024; Tjuatja et al., 2024). Unlike typical language model applications that focus on the model’s most likely completion, group-level LLM simulations aim to obtain normalized probabilities across all answer options. Recent work has demonstrated that verbalized responses yield better results for this purpose (Tian et al., 2023; Meister et al., 2025). Nevertheless, calibration of LLM outputs remains an open challenge; while extensively studied for model answer confidence (Zhao et al., 2021; Jiang et al., 2021; Kapoor et al., 2024; Zhu et al., 2023) and hallucinations (Kalai & Vempala, 2024), these issues also apply to simulating population response distributions. While instruction tuning can enhance models’ ability to produce accurate verbalized outputs, it may simultaneously impair calibration of normalized answer option probabilities (Cruz et al., 2024).

### v4

@@ 6 Conclusion and Future Work
Using SimBench, we provide the first systematic analysis of simulation ability across 45 LLMs. We show that even the best LLMs have limited simulation ability, that performance scales log-linearly with model size, and that there exists a fundamental tradeoff between standard alignment and simulating diverse human opinions. We further show that models struggle more when simulating specific demographic groups, and that strong simulation ability correlates with deep, knowledge-intensive reasoning. While significant progress is needed, SimBench makes this progress measurable, providing an open foundation to accelerate the development of more faithful LLM simulators.
We hope that future research can build on this foundation by expanding beyond standardized formats to evaluate interactive and open-ended behavioral simulations. Finally, addressing the observed tradeoff between alignment and simulation fidelity requires developing distribution-preserving alignment techniques, alongside further investigation into the causal mechanisms linking other model capabilities to simulation performance.
Acknowledgments
This work was partially conducted while TH, JB, and PR were at the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis. TH was supported by the Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation). JB acknowledges funding from the Swiss National Science Foundation (SNSF) through project P500-2_235328. DH was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 949944, INTEGRATOR). PR was supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA).
Ethics Statement
SimBench’s primary purpose is to benchmark LLM ability to simulate human behavior. While advancements in LLM simulation capabilities can support helpful applications such as pre-testing policies, these do not come without risks of misrepresentation and dual use.
Responsible Use and Acknowledgment of Limitations
@@ Acknowledgments
Data Provenance and Transformative Use
@@
Scope of Representation and Intersectional Analysis
@@ Conclusion
Reproducibility Statement
SimBench is released on HuggingFace. Our codebase is available on GitHub. We provide detailed descriptions of our experimental setup, including the exact prompts used for both base and instruction-tuned models, in Appendix D, and an empirical validation of our elicitation methodology in Appendix E.
@@ Appendix A Limitations
The challenge of benchmarking human behavior lies in its vastness. No single study, nor dozens, could ever capture its full complexity. We operationalize this challenge by focusing on a core set of fundamental cognitive and social tasks widely studied in the behavioral sciences: decision-making, self-assessment, judgment, and problem-solving. This focus necessarily means that SimBench does not capture the full complexity of human interaction, such as embodied or multi-turn social dynamics. As the first large-scale benchmark for group-level simulation, SimBench provides the essential infrastructure to establish robust baselines, uncover fundamental properties like scaling trends, and map the frontier for future work on more complex, interactive simulations.
@@ Appendix C SimBenchPop and SimBenchGrouped Sampling Details
For SimBenchGrouped, we focus only on five large-scale survey datasets with rich demographic information and sufficient sample sizes: OpinionQA, ESS, Afrobarometer, ISSP, and Latinobarómetro. Our sampling approach prioritizes questions showing meaningful demographic variation. For each dataset, we identify available grouping variables (e.g., age, gender, country) with sufficient group sizes to form meaningful response distributions. We calculate the variance of responses across demographic groups for each question and rank questions by their variance scores, prioritizing those showing the strongest demographic differences. We select questions that exhibit significant variation across demographic groups to ensure the benchmark captures meaningful differences in responses. For each selected question, we create multiple test cases by pairing it with different values of the grouping variables (e.g., age = “18-29”, age = “30-49”). This process results in 6,343 test cases that specifically measure LLMs’ ability to simulate responses from narrower participant groups based on specified demographic characteristics. Table 4 provides a summary of the sampling process across all datasets, showing the minimum group size thresholds and the number of test cases in each benchmark split.
Table 4: Dataset Sampling Summary; NaN refers to a dataset that is only available in aggregated form and no grouping size is known.
@@ Appendix E Validation of Elicitation Method
Figure 5 compares the SimBench scores for several instruction-tuned models using both methods. The results are unequivocal: using verbalized distributions (teal dots) dramatically and consistently outperforms direct token probabilities (orange dots) for every instruction-tuned model tested. In many cases, using token probabilities results in scores far below zero, indicating that the model’s raw logits are poorly calibrated for this task after instruction tuning. In contrast, base models (black bars) perform reasonably well with token probabilities, as they are not subject to the same post-training shifts.
@@ Appendix G Full SimBench Results (RQ1)
We show the SimBench scores for all 45 models we evaluate in Table 7. We show the scaling law plots for all models in Figure 6.
@@ H.2 Within-Family Performance Analysis
To investigate the impact of model size (RQ2), we conducted one-way ANOVA followed by post-hoc Tukey HSD tests for pairwise comparisons within model families. The results, presented in Table 9, show that for most families, increases in model size lead to statistically significant improvements in simulation ability, supporting the scaling trends observed in Figure 2.
@@
C(dataset_name)[T.Latinobarómetro] -0.0368 0.0019 -18.9530 0.0000 -0.0406 -0.0330
@@ Conclusion
These results provide strong evidence for inconsistent mediation and resolve a key paradox in our findings. While our analysis in §4.4 shows that instruction tuning harms simulation fidelity on high-entropy questions, our main leaderboard (Table 1) shows that the best overall simulators are instruction-tuned. This mediation analysis explains why: the total effect of instruction tuning is the net outcome of two larger, opposing forces. First, a direct positive effect (+6.46 points) on capability, likely from improved instruction- and persona-following. Second, a smaller but significant indirect negative effect (-1.74 points) caused by entropy suppression. The net positive effect (+4.72 points) demonstrates that, on average, the direct benefits of alignment currently outweigh the harm from reduced distributional diversity. Future work on creating SOTA simulators should therefore focus on developing hybrid or “distribution-preserving” alignment methods that retain the direct benefits of instruction tuning while mitigating its harmful, entropy-reducing side effects.
@@
This qualitative divergence is key. The two methods are not just different in degree, but in kind. Instruction tuning boosts performance by sharpening a model’s ability to follow prompt instructions and converge on a consensus answer. Specialist tuning boosts performance by aligning the model’s internal representations more closely with the patterns of human choice. Because they target different mechanisms, their benefits are not mutually exclusive. This suggests that perhaps future gains in LLM simulation will come from hybrid approaches that synthesize both paradigms, creating models that are both generally capable and foundationally aligned with the nuances of human behavior.
@@
Questions: 114, with an average of 518 responses per question.
@@
Participant grouping variables: None.
Default System Prompt:
@@
Participants: roughly 10,000 paid participants from a representative sample of the US populace.
@@
Description: This dataset contains three moral decision-making scenarios, for which participants provided subjective choices. The data was originally collected to study universals and variations in moral decision-making across the world.
@@
Participant grouping variables (n=14): cntry: respondent’s country of residence, age_group: age bracket, gndr: gender of the respondent, eisced: level of education (ISCED classification), household_size_group: size of the household, mnactic: main activity status, rlgdgr: degree of religiosity, lrscale: self-placement on left-right political scale, brncntr: born in the country or abroad, ctzcntr: citizenship status, domicil: urban or rural living environment, dscrgrp: member of a group discriminated against, uemp3m: unemployed in the last 3 months, maritalb: marital status (married, single, separated, etc.).
@@
Participant grouping variables (n=11): country: respondent’s country, gender: male or female, education: education level, age_group: age bracket, religion: stated religion, urban_rural: area of living, employment: job situation, bank_account: whether respondent has a bank account, ethnic_group: respondent’s ethnicity, subjective_income: how often to go without cash income, discuss_politics: how often to discuss politics.
@@
Participant grouping variables (n=22): age_group: age bracket, gender_cat: male or female or other, race_cat: respondent’s race, engnat_cat: is (not) English native, hand_cat: right/left/both-handed, education_cat: level of education, urban_cat: type of urban area, religion_cat: stated religion, orientation_cat: sexual orientation, voted: did (not) vote at last elections, married: never/currently/previously, familysize: number of people belonging to the family, TIPI_E_Group: extraversion level based on TIPI score, TIPI_A_Group: agreeableness level based on TIPI score, TIPI_C_Group: conscientiousness level based on TIPI score, TIPI_ES_Group: emotional stability level based on TIPI score, TIPI_O_Group: openness-to-experience level based on TIPI score, household_income: income sufficiency, work_status: job situation, nr_of_persons_in_household: 1-7+, marital_status: respondent’s legal relationship status, domicil: type of urban area.
Default System Prompt:
@@
Participant grouping variables (n=11): country: respondent’s country, age: age bracket, gender: male or female, years_of_education: 1-10+, household_income: income sufficiency, work_status: job situation, religion: stated religion, nr_of_persons_in_household: 1-7+, marital_status: respondent’s legal relationship status, domicil: type of urban area, topbot: self-assessed social class.
@@
Looking at the list below, who do you think should have the greatest responsibility for reducing differences in income between people with high incomes and people with low incomes? Options: (A): Can’t choose (B): Private companies (C): Government (D): Trade unions (E): High-income individuals themselves (F): Low-income individuals themselves (G): Income differences do not need to be reduced
@@
Participants: In total, 19,205 interviews were applied in 18 countries. Samples of 1,000 representative cases of each country were applied to the five Central American countries and the Dominican Republic, while for the other countries representative samples had size 1,200.
@@
Participant grouping variables (n=1): country.
@@
Description: This dataset comprises 990 multi-turn adversarial conversations between humans and a dialogue model, accompanied by raters’ judgment on the safety of the last model response. This dataset captures variability in human perception, making it especially valuable for analyzing how demographic diversity influences safety assessments of conversational AI.
Questions: 27,748 (990 turns annotated in multiple dimensions), with an average of 72 responses per question.
@@
Participant grouping variables (n=4): Country: country of origin, Age_Group: age bracket of the respondent, Gender: gender of the respondent, Education: highest level of education completed.
Default System Prompt:
@@
Participant grouping variables (n=1): UserCountry3: participant country.
Default System Prompt:
@@
Description: This dataset includes judgments about individuals’ perception of science, its role in society and politics, attitudes toward climate change, and science communication behaviors.
Questions: 97, with an average of 69,234 responses per question.
@@
Prior research has primarily relied on first token probabilities to obtain survey answers from LLMs (Santurkar et al., 2023; Dominguez-Olmedo et al., 2024; Tjuatja et al., 2024). Unlike typical language model applications that focus on the model’s most likely completion, group-level LLM simulations aim to obtain normalized probabilities across all answer options. Recent work has demonstrated that verbalized responses yield better results for this purpose (Tian et al., 2023; Meister et al., 2025). Nevertheless, calibration of LLM outputs remains an open challenge; while extensively studied for model answer confidence (Zhao et al., 2021; Jiang et al., 2021; Kapoor et al., 2024; Zhu et al., 2023) and hallucinations (Kalai & Vempala, 2024), these issues also apply to simulating population response distributions Hu et al. (2026). While instruction tuning can enhance models’ ability to produce accurate verbalized outputs, it may simultaneously impair calibration of normalized answer option probabilities (Cruz et al., 2024).
@@
Appendix O Detailed Demographic Breakdown
Table LABEL:tab:country_samples provides a detailed count of the number of simulation targets (samples) associated with specific countries or regions across the SimBenchPop and SimBenchGrouped splits.
Table 13: Distribution of samples by Country/Region. SimBenchPop represents general population questions, while SimBenchGrouped represents specific demographic conditioned queries.
Albania 0 5 5
Angola 12 11 23
Argentina 37 51 88
Armenia 0 2 2
Australia 6 188 194
Austria 72 52 124
Bangladesh 0 20 20
Belarus 0 1 1
Belgium 90 39 129
Benin 39 10 49
Bolivia 120 33 153
Botswana 52 24 76
Brazil 72 104 176
Bulgaria 78 20 98
Burkina Faso 45 11 56
Cabo Verde 14 7 21
Cameroon 90 21 111
Canada 0 203 203
Chile 66 51 117
China 9 38 47
Colombia 105 33 138
Costa Rica 35 37 72
Côte d’Ivoire 75 28 103
Croatia 12 27 39
Cyprus 10 8 18
Czech Republic 108 49 157
Democratic Republic of the Congo 0 3 3
Denmark 60 42 102
Dominican Republic 90 33 123
Ecuador 79 36 115
Egypt 0 16 16
El Salvador 36 25 61
Estonia 40 26 66
Eswatini 39 9 48
Ethiopia 80 26 106
Finland 64 54 118
France 149 132 281
Gabon 30 11 41
Gambia 42 14 56
Georgia 10 11 21
Germany 146 236 382
Ghana 50 20 70
Greece 0 28 28
Guatemala 88 33 121
Guinea 28 17 45
Honduras 57 31 88
Hong Kong 0 10 10
Hungary 92 53 145
Iceland 36 45 81
India 7 343 350
Indonesia 0 16 16
Iran 0 2 2
Iraq 0 1 1
Ireland 102 35 137
Israel 70 39 109
Italy 105 70 175
Japan 52 36 88
Jordan 0 10 10
Kazakhstan 0 12 12
Kenya 100 34 134
Kuwait 0 3 3
Latvia 0 2 2
Lebanon 0 12 12
Lesotho 28 16 44
Liberia 12 13 25
Lithuania 113 43 156
Madagascar 42 12 54
Malawi 48 8 56
Malaysia 0 18 18
Mali 14 16 30
Mauritania 52 14 66
Mauritius 12 12 24
Mexico 65 63 128
Mongolia 0 1 1
Montenegro 20 5 25
Morocco 13 32 45
Mozambique 33 12 45
Myanmar 0 2 2
Namibia 45 10 55
Netherlands 35 41 76
New Zealand 18 37 55
Nicaragua 0 9 9
Niger 52 16 68
Nigeria 19 41 60
North Macedonia 0 10 10
Norway 25 31 56
Pakistan 0 15 15
Palestinian Territory 0 5 5
Panama 45 29 74
Paraguay 43 24 67
Peru 119 39 158
Philippines 37 70 107
Poland 111 61 172
Portugal 80 28 108
Puerto Rico 0 2 2
Republic of the Congo 30 11 41
Romania 0 12 12
Russia 159 85 244
São Tomé and Príncipe 26 8 34
Senegal 42 14 56
Serbia 20 9 29
Seychelles 14 19 33
Sierra Leone 45 9 54
Singapore 0 4 4
Slovakia 50 29 79
Slovenia 89 42 131
South Africa 54 53 107
South Korea 5 24 29
Spain 61 63 124
Sudan 13 5 18
Suriname 10 15 25
Sweden 71 52 123
Switzerland 132 38 170
Taiwan 26 21 47
Tajikistan 0 1 1
Tanzania 46 19 65
Thailand 48 20 68
Togo 30 11 41
Tunisia 11 18 29
Türkiye 24 30 54
Uganda 48 22 70
Ukraine 0 11 11
United Kingdom 74 263 337
United States 1028 1636 2664
Uruguay 54 42 96
Uzbekistan 0 2 2
Venezuela 147 45 192
Vietnam 0 8 8
Zambia 64 16 80
Zimbabwe 42 18 60
Grand Total 6,009 6,343 12,352